Show that T:R^3 →R^2: T(x,y,z)= (2x+y-z,x+z) is a linear transformation. Verify that T satisfies the Rank-nullity theorem


1
Expert's answer
2021-03-03T04:21:48-0500

Let us show that "T: \mathbb R^3 \to \mathbb R^2,\ \ T(x,y,z) = (2x+y-z,x+z)," is a linear transformation.

Let "a,b\in \mathbb R,\ (x_1,y_1,z_1),(x_2,y_2,z_2)\in\mathbb R^3." Then "T[a(x_1,y_1,z_1)+b(x_2,y_2,z_2)]=T[(ax_1+bx_2, ay_1+by_2, az_1+bz_2)]=\n(2(ax_1+bx_2)+ay_1+by_2-(az_1+bz_2), ax_1+bx_2+az_1+bz_2)=\n(2ax_1+2bx_2+ay_1+by_2-az_1-bz_2, ax_1+bx_2+az_1+bz_2)=\n(a(2x_1+y_1-z_1)+b(2x_2+y_2-z_2), a(x_1+z_1)+b(x_2+z_2))=\n(a(2x_1+y_1-z_1), a(x_1+z_1))+(b(2x_2+y_2-z_2), b(x_2+z_2))=\na(2x_1+y_1-z_1, x_1+z_1)+b(2x_2+y_2-z_2, x_2+z_2)=\naT(x_1,y_1,z_1)+bT(x_2,y_2,z_2)"

Therefore, "T" is a linear transformation.


Let us find its kernel:


"\ker(T)=\{(x,y,z)\in\mathbb R^3\ |\ T(x,y,z)=(0,0)\}=
\{(x,y,z)\in\mathbb R^3\ |\ (2x+y-z,x+z)=(0,0)\}=
\{(x,y,z)\in\mathbb R^3\ |\ 2x+y-z=0, x+z=0\}=
\{(x,y,z)\in\mathbb R^3\ |\ y=z-2x, x=-z\}=
\{(x,y,z)\in\mathbb R^3\ |\ y=3z, x=-z\}=
\{(-z,3z,z)\ |\ z\in\mathbb R\}=\{z(-1,3,1)\ |\ z\in\mathbb R\}"


It follows that "\dim(\ker T)=1."


Let us find "Im(T)." Let "(a,b)\in\mathbb R^2" be arbitrary. Then "T[(b, a-2b, 0)]=(2b+a-2b+0, b+0)=(a,b)", and therefore, "T" is surjective. It followws that "\dim(Im T)=\dim(\mathbb R^2)=2"


Consequently, "T" satisfies the Rank-nullity theorem:


"\dim(\ker T)+\dim(Im T)=1+2=3=\dim(\mathbb R^3)."